The class average on a test was 81, with a standard deviation of 2.6. Find the probability that a student received at least a 78 on the test.
Answer by stanbon(75887)

The class average on a test was 81, with a standard deviation of 2.6. Find the probability that a student received at least a 78 on the test.
-------------------
z(78) = (78-81)/2.6 = -3/2.6 = -1.1538
-----
P(x >= 78) = P(z >= -1.1538) = normalcdf(-1.1538,100) = 0.8757
===============
Cheers,
Stan H.
===============